OpenStudy (emily_98):

Aisha wants to make two quilts, each with the same area. The first quilt will be square with sides s feet long. The second quilt will be a rectangle with a width that is half the length of a side of the square quilt and a length that is 6 feet longer than a side length of the square quilt. Which quadratic equation can be used to find s, the side length of the square quilt? s2 = (s + 6) s2 = (s)(s + 6) s2 = (6s) s2 = (s)(6s)

OpenStudy (benlindquist):

@netflixandno

OpenStudy (isaidavila):

First find the sides of the rectangle Width is s/2 Lengt s+6 Both áreas are equal A1=A2 S^2=(s/2)*(s+6) 2s^2=s^2+6s S^2=6s

umerlodhi (umerlodhi):